OpenStudy (anonymous):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y A. There is no solution. B. There are an infinite number of solutions. C. The solution is (1, 10) D. The solution is (28, 0) There's no solution for this right?

OpenStudy (anonymous):

substitution method refers to a method in which value of a variable is substituted from one equation to the other one!

OpenStudy (anonymous):

y – 3x = 7 21 + 9x = 3y from first equation we get the ba;ue of y = 7 +3x now using it in second one we get 21 + 9x = 3 ( 7 + 3x) => 21 + 9x = 21 + 9x => 0 = 0 there is no solution for it for it you are correct :D

OpenStudy (anonymous):

ok sorry '.'

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

I thought I was wrong XD Thank you :)

OpenStudy (anonymous):

your welcome :D

OpenStudy (anonymous):